Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, December 21, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 21, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/21/98 (Monday)

Segment 1 report for 12/21/98: Due to snow and icy conditions, I-5 was closed from Templin Hwy. to the Grapevine until ~10:00. Crews were unable to work in sites of construction in that area; Cuddy Creek Tributary, Lebec Road (at the Warbler skip), and on Tejon Bench. On Grapevine Rd., backfill continued to 15.85. Installation of casing continued at the Stream Crossing #9 (EMP 15.35) and the Tejon Bench/I-5 (EMP 16.49) bores. Welding, x-ray and coating continued on the south side of the California Aqueduct bore (EMP 12.28). Backfill also began there and was in progress at the Stream Crossing #4 bore (EMP 9.03) and the Tecuya Creek/I-5 bore (EMP 9.86). Electrical work continued at the Grapevine Pump Station. Testing of the equipment was in progress at the Emidio Pump Station.

Segment 2B (Castaic to Sylmar): Report for 12-21-98. Crews worked to complete paving at several sites along the right of way. Areas included Lyons Ave, The Old Rd south of Chiquella Ln and The Old Rd south of Magic Mtn Pkwy. In addition, crews were locating test leads along the right of way that had been covered with paving. Finally, clean up crews continued work at Weldon Cyn removing spoils and debris.

Segment 3B (Sylmar to Los Angeles): Report for 12-21-98. At Hubbard St crews progressed with clean up of the rail right of way while crews worked on the pipe at the Trumman St bridge. Near Tuxford St crews set a hand box for fiber optic duct and backfilled the bore pit. Near Hollywood Way, fiber optic cable was pulled while crews working at Cornfield yard installed a muffler on the end of the pipe for hydro testing and began fabrication of the de-watering pipe.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page